if Na+ are placed at fcc lattice point and Cl- are placed at all O.H.V. then what will be the structure with respect to Cl-

Grade:12th pass

1 Answers

Vikas TU
7 years ago
The contribution of Na+ is:
Na+ = 8*1/8 +  1*6/2 = > 3 + 1 =>  4
The contribution of Cl- is:
Cl- = 1*6/2 = 3
Therefore the formulae is=>
Na4Cl3